# A determinantal formula for orthosymplectic Schur functions

### v1

@@ 1. Introduction
Hook Schur polynomials (or supersymmetric Schur polynomials) and orthosymplectic Schur polynomials are the characters of irreducible representations of general linear and orthosymplectic Lie superalgebras respectively. These are supersymmetric functions in two sets of commuting variables indexed by integer partitions. It is well known that hook Schur polynomials form the basis for the ring of supersymmetric functions. For background on the representation theory of Lie superalgebras, see [CW12].
Hook Schur polynomials were introduced by Berele and Regev [BR87] combinatorially using hybrid tableaux involving a column-strict part and a row-strict part. Stembridge [Ste85] gave a different characterization for these polynomials and showed that they form a basis for the ring of supersymmetric functions. Moens and Jeugt [MvdJ03] proved the following determinantal formula for these polynomials using the representation theory of general linear superalgebra $\mathfrak{gl}(m/n)$.
@@
Let $\lambda$ be a partition such that $\lambda_{n+1}\leq m$ and $k=\min\{i|\lambda_{i}+n+1-i\leq m\}.$ Then the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is given by
$\operatorname{hs}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}\left(\frac{1}{x_{i}+y_{j}}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&\left(x_{i}^{\lambda_{j}+n-m-j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ \\ \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(1.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right)$
@@
$\operatorname{spo}_{\lambda}(x_{1}/y_{1},y_{2})=\frac{(x_{1}+y_{1})(x_{1}+y_{2})(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}{(x_{1}-\bar{x}_{1})(y_{1}-y_{2})}\\ \times\left(\frac{x_{1}}{(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}-\frac{\bar{x}_{1}}{(x_{1}+y_{1})(x_{1}+y_{2})}\right)\times\det\left(\begin{array}[]{cc}y_{1}&y_{2}\\ 1&1\end{array}\right).$
@@
$x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
@@
$\operatorname{spo}_{\lambda}(x_{1},x_{2},\dots,x_{n}/y)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
@@
We prove 1.2 in Section 3 after giving all required definitions in Section 2. Our proof uses the Jacobi–Trudi type formula for orthosymplectic Schur polynomials. The Jacobi–Trudi type formula was given by Balantekin and Bars [BB81]. It was proved using a different approach by Benkart, Shader and Ram [BSR98] and recently, Stokke and Visentin [SV16] constructed the formula using lattice paths. Orthosymplectic Cauchy identities and Pieri rule for orthosymplectic Schur polynomials were proved in [PPS22] and  [Sto18] respectively.
We also show that the odd symplectic characters introduced by Proctor [Pro88] are the same as the orthosymplectic Schur polynomials with some specialized indeterminates (see 2.11). Lastly, we give a generalization of an odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals (see  [BKW16, BKW15, Oka20]) in Section 4.
@@ 2. Preliminaries
Recall that a partition $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{\ell})$ is a sequence of weakly decreasing nonnegative integers. Any entry of the partition is called a part. The length of a partition $\lambda$, which is the number of positive parts, is denoted by $\ell(\lambda)$. A partition $\lambda$ can be represented pictorially as a Young diagram, whose $i$’th row contains left-justified boxes. We will use the so-called English notation where the first row is on top. The conjugate of a partition $\lambda$, denoted $\lambda^{\prime}$, is the partition whose Young diagram is obtained by transposing the Young diagram of $\lambda$.
@@
(2.3) $s_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j}\Bigr{)}}.$
The expression in (2.3) is essentially the Weyl character formula. The following Jacobi–Trudi identity expresses the Schur polynomial $s_{\lambda}(X)$ in terms of complete symmetric functions,
(2.4) $s_{\lambda}(X)=\det\left(h_{\lambda_{i}-i+j}(X)\right)_{1\leq i,j\leq r},$
where $r$ is any integer such that $r\geq\ell(\lambda)$. Recall that there is a combinatorial formula for $s_{\lambda}(X)$ as a sum of monomials summed over all column-strict Young tableaux of shape $\lambda$. A column-strict Young tableau of shape $\lambda$ is a filling of $\lambda$ with entries from the set $[n]$ such that
  • the entries increase weakly along rows
  • the entries increase strictly along columns.
For partitions $\lambda$ and $\mu$ such that $\mu\subset\lambda\,(\mu_{i}\leq\lambda_{i}$, for all $i\geq 1)$, the skew shape $\lambda/\mu$ is the set theoretic difference $\lambda/\mu$ and the skew Schur function $s_{\lambda/\mu}(X)$ is given by
(2.5) $s_{\lambda/\mu}(X)=\sum_{T}\operatorname{wt}(T),\qquad\operatorname{wt}(T)\coloneqq\prod_{i\geq 1}x_{i}^{c_{i}(T)},$
where the sum is taken over all column-strict tableaux of shape $\lambda/\mu$ and $c_{i}(T)$, $i\in[n]$ is the number of occurrences of $i$ in $T$. Otherwise $s_{\lambda/\mu}(X)=0$. The following Jacobi–Trudi formula expresses $s_{\lambda/\mu}(X)$ in terms of complete symmetric functions.
(2.6) $s_{\lambda/\mu}(X)=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(X)\right)_{1\leq i,j\leq r},$
@@
We now give a supersymmetric analogue of the symmetric functions. We consider the ring $\mathbb{Z}[x_{1},\dots,x_{n},y_{1},\dots,y_{m}]$ of polynomials in $(n+m)$ commuting variables $x_{1},\dots,x_{n},$ $y_{1},\dots,y_{m}$ with integer coefficients. A polynomial $f(X,Y)$ in this ring is doubly symmetric if it is symmetric in both $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{m})$. Moreover, if substituting $x_{n}=t$ and $y_{m}=-t$ results in an expression independent of $t$, then we call $f(X,Y)$ a supersymmetric function. The elementary supersymmetric function and the complete supersymmetric function indexed by $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ are:
(2.7) $E_{\lambda}(X/Y)=\prod_{i=1}^{n}E_{\lambda_{i}}(X/Y),$
@@
$E_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}e_{j}(X)h_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0,\end{cases}$
@@
(2.8) $H_{\lambda}(X/Y)=\prod_{i=1}^{n}H_{\lambda_{i}}(X/Y),$
@@
$H_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}h_{j}(X)e_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0.\end{cases}$
It is convenient to define $E_{r}(X/Y)$ and $H_{r}(X/Y)$ to be zero for $r<0$.
@@
  • the entries increase weakly along rows and columns
  • the unprimed entries strictly increase along columns
  • the primed entries strictly increase along rows
@@
$\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{n_{i}(T)}\prod_{j=1^{\prime}}^{m^{\prime}}y_{j}^{n_{j}(T)},$
where $n_{\alpha}(T)$, for $\alpha$ being any of the entries given in 2.1, is the number of occurrences of $\alpha$ in $T$. Denote the set of supertableaux of skew shape $\lambda$ with entries given in 2.1 by $\operatorname{SSYT}_{n/m}(\lambda)$. For an integer partition $\lambda$, the hook Schur (or supersymmetric Schur) polynomial, denoted $\operatorname{hs}_{\lambda}(X/Y)$ is given by
(2.9) $\operatorname{hs}_{\lambda/\mu}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{SSYT}_{n/m}(\lambda)}\operatorname{wt}(T).$
@@
$\operatorname{hs}_{(2,1)}(x_{1},x_{2}/y_{1})=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}y_{1}+2x_{1}x_{2}y_{1}+x_{2}^{2}y_{1}+x_{1}y_{1}^{2}+x_{2}y_{1}^{2}.$
The Jacobi–Trudi formula expresses $\operatorname{hs}_{\lambda}(X/Y)$ in terms of the complete supersymmetric functions [PT92].
(2.10) $\operatorname{hs}_{\lambda}(X/Y)=\det\left(H_{\lambda_{i}-i+j}(X/Y)\right)_{1\leq i,j\leq n}.$
Note that the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. Now we consider orthosymplectic Schur functions after defining symplectic Schur functions.
@@
A symplectic tableau $T$ of shape $\lambda$ with entries
$1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}$
  • the entries increase weakly along rows and strictly along columns
[…]
@@
$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)},$
where $n_{i}(T)$ and $n_{\overline{i}}(T)$ is the number of occurrences of $i$ and $\overline{i}$ in $T$ respectively. Denote the set of symplectic tableaux of shape $\lambda$ filled with entries given in 2.3 by $\operatorname{SPT}_{n}(\lambda)$.
@@
The following is a symplectic tableau $T$ of shape $\lambda=(3,2,2)$ with $\operatorname{wt}(T)=x_{2}$.
@@
$\mathrm{sp}_{\lambda}(X)\coloneqq\sum_{T\in\operatorname{SPT}_{n}(\lambda)}\operatorname{wt}(T),$
where the sum runs over the set of tableaux in $\operatorname{SPT}_{n}(\lambda)$. The symplectic Schur function can be expressed using the Weyl character formula in the following way. Let $\lambda$ be a partition of length at most $n$. Then
(2.11) $\mathrm{sp}_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}},$
@@
(2.12) $\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\,\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
@@
An orthosymplectic tableau $T$ of shape $\lambda$ is a filling of Young diagram of $\lambda$ with entries from $B$ satisfying the following conditions:
  • the portion $S$ of $T$ that contains entries only from $B_{0}$ is symplectic.
@@
$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)}\prod_{j=1}^{m}y_{j}^{n_{i^{\prime}}(T)},$
where $n_{\alpha}$ is the number of occurrences of $\alpha$, for $\alpha$ being any of the entries given in 2.5. Denote the set of orthosymplectic tableaux of shape $\lambda$ filled with entries $1\dots m^{\prime}$ by $\operatorname{OSPT}_{n/m}({\lambda})$.
@@
The following is an orthosymplectic tableau $T$ of shape $\lambda=(3,3,2,1)$ and $\operatorname{wt}(T)=x_{1}x_{4}^{-1}y_{1}y_{2}y_{3}^{3}$.
@@
$\operatorname{spo}_{\lambda}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{OSPT}_{n/m}({\lambda})}\operatorname{wt}(T).$
Note that $\operatorname{spo}_{\lambda}(X/Y)$ is zero unless $\lambda_{n+1}\leq m$. Also, $\operatorname{spo}_{\lambda}(X/\emptyset)=\mathrm{sp}_{\lambda}(X)$. Moreover,
$\operatorname{spo}_{\lambda}(X/Y)=\sum_{\mu}\mathrm{sp}_{\mu}(X)s_{\lambda^{\prime}/\mu^{\prime}}(Y).$
@@
Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then we have the following orthosymplectic tableaux in
@@
(2.13) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{n\times n},$
where $J_{r}=\sum_{l=0}^{r}h_{l}(X,\bar{X})e_{r-l}(Y)$.
Note that the orthosymplectic Schur polynomial $\operatorname{spo}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. The following remark gives an expression for $\operatorname{spo}_{\lambda}(X/y)$ if $\lambda_{n+1}\leq 1$.
@@
[…]
$\operatorname{spo}_{\lambda}(X/y)=\frac{y^{\ell-n}}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})\right),$
@@
Let $\lambda$ be a partition of length at most $n$. A $(n-1,1)$-symplectic tableau of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from
@@
  • the entries in each row are weakly increasing;
  • the entries in each column are strictly increasing;
  • the entries of the $i$’th row are greater than or equal to $i$
We denote by $\operatorname{SPT}^{(n-1,1)}(\lambda)$ the set of $(n-1,1)$-symplectic tableaux of shape $\lambda$. Given a $(n-1,1)$-symplectic tableau $T$, we define
@@
where $\alpha_{i}(T)$, $\alpha_{\overline{i}}(T)$ for $i\in[n-1]$, and $\alpha_{n}(T)$ are number of occurences of $i$, $\overline{i}$ and $n$ respectively.

### v2

@@ 1. Introduction
Hook Schur polynomials (or supersymmetric Schur polynomials) and orthosymplectic Schur polynomials are the characters of irreducible representations of general linear and orthosymplectic Lie superalgebras respectively. These are supersymmetric functions in two sets of commuting variables indexed by integer partitions. It is well known that hook Schur polynomials form a basis for the ring of supersymmetric functions. For background on the representation theory of Lie superalgebras, see [CW12].
Hook Schur polynomials were introduced by Berele and Regev [BR87] combinatorially using hybrid tableaux involving a column-strict part and a row-strict part. Stembridge [Ste85] gave a different characterization for these polynomials and showed that they form a basis for the ring of supersymmetric functions. Moens and Jeugt [MvdJ03] proved the following determinantal formula for these polynomials using the representation theory of the general linear superalgebra $\mathfrak{gl}(m/n)$.
@@
Let $\lambda$ be a partition such that $\lambda_{n+1}\leq m$ and $k=\min\{i\mid\lambda_{i}+n+1-i\leq m\}.$ The hook Schur polynomial $\operatorname{hs}_{\lambda}(X;Y)$ is given by
$\operatorname{hs}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}\left(\frac{1}{x_{i}+y_{j}}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&\left(x_{i}^{\lambda_{j}+n-m-j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ \\ \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j\mid\lambda_{j}+n+1-j\leq m\}.$ The orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
(1.1) $\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right)$
@@
$\operatorname{spo}_{\lambda}(x_{1};y_{1},y_{2})=\frac{(x_{1}+y_{1})(x_{1}+y_{2})(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}{(x_{1}-\bar{x}_{1})(y_{1}-y_{2})}\\ \times\left(\frac{x_{1}}{(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}-\frac{\bar{x}_{1}}{(x_{1}+y_{1})(x_{1}+y_{2})}\right)\times\det\left(\begin{array}[]{cc}y_{1}&y_{2}\\ 1&1\end{array}\right).$
@@
(1.2) $x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
As the expression for $\operatorname{spo}_{(2)}(x_{1};y_{1},y_{2})$ in (1.2) is the same as the one given in (2.12), 1.2 is verified in this instance.
@@
$\operatorname{spo}_{\lambda}(x_{1},x_{2},\dots,x_{n};y)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
@@
We prove 1.2 in Section 3 after giving all required definitions in Section 2. Our proof uses the Jacobi–Trudi type formula for orthosymplectic Schur polynomials. The Jacobi–Trudi type formula was given by Balantekin and Bars [BB81]. It was proved using a different approach by Benkart, Shader and Ram [BSR98] and recently, Stokke and Visentin [SV16] proved the formula using lattice paths. Orthosymplectic Cauchy identities and a Pieri rule for orthosymplectic Schur polynomials were proved in [PPS22] and  [Sto18] respectively.
We also show that the odd symplectic characters introduced by Proctor [Pro88] are the same as the orthosymplectic Schur polynomials with some specialized indeterminates (see 2.12). Lastly, we give a generalization of an odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals (see  [BKW16, BKW15, Oka20]) in Section 4.
@@ 2. Preliminaries
Recall that a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$ is a sequence of weakly decreasing nonnegative integers containing only finitely many non-zero terms. The length of a partition $\lambda$, which is the number of positive parts, is denoted by $\ell(\lambda)$. A partition $\lambda$ can be represented pictorially as a Young diagram, whose $i^{\text{th}}$ row contains $\lambda_{i}$ left-justified boxes. We will use the so-called English notation where the first row is on top. The conjugate of a partition $\lambda$, denoted $\lambda^{\prime}$, is the partition whose Young diagram is obtained by transposing the Young diagram of $\lambda$. For example, Figure 1 shows the Young diagrams of $\lambda=(3,2,2)$ and its conjugate $\lambda^{\prime}=(3,3,1)$.
[…]
3,2,2
[…]
3,3,1
Figure 1. The Young diagram of $\lambda=(3,2,2)$ and $\lambda^{\prime}=(3,3,1)$
@@
(2.3) $\mathrm{s}_{\lambda}(X)=\frac{\displaystyle\det\Bigl{(}x_{i}^{\lambda_{j}+n-j}\Bigr{)}_{1\leq i,j\leq n}}{\displaystyle\det\Bigl{(}x_{i}^{n-j}\Bigr{)}_{1\leq i,j\leq n}}.$
The expression in (2.3) is essentially the Weyl character formula.
For partitions $\lambda$ and $\mu$ such that $\mu\subset\lambda\,(\mu_{i}\leq\lambda_{i}$, for all $i\geq 1)$, the skew shape $\lambda/\mu$ is the set theoretic difference where $\lambda$ and $\mu$ are regarded as Young diagrams, i.e., collections of boxes. Recall that a semistandard Young tableau of shape $\lambda/\mu$ is a filling of $\lambda/\mu$ such that
  • the entries increase weakly across rows,
  • the entries increase strictly down columns.
The skew Schur function $\mathrm{s}_{\lambda/\mu}(X)$ is given by
(2.4) $\mathrm{s}_{\lambda/\mu}(X)=\sum_{T}\operatorname{wt}(T),\qquad\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{c_{i}(T)},$
where the sum is taken over all semistandard Young tableaux of shape $\lambda/\mu$ with entries from the set $[n]$ and $c_{i}(T)$, $i\in[n]$ is the number of occurrences of $i$ in $T$. If $\mu=\emptyset$, then $\mathrm{s}_{\lambda/\mu}(X)=\mathrm{s}_{\lambda}(X)$. The following Jacobi–Trudi formula expresses $\mathrm{s}_{\lambda/\mu}(X)$ in terms of complete symmetric functions:
(2.5) $\mathrm{s}_{\lambda/\mu}(X)=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(X)\right)_{1\leq i,j\leq r},$
@@
We now give a supersymmetric analogue of the symmetric functions. We consider the ring $\mathbb{Z}[x_{1},\dots,x_{n},y_{1},\dots,y_{m}]$ of polynomials in $(n+m)$ commuting variables $x_{1},\dots,x_{n},$ $y_{1},\dots,y_{m}$ with integer coefficients. A polynomial $f(X,Y)$ in this ring is doubly symmetric if it is symmetric in both $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{m})$. Moreover, if substituting $x_{n}=t$ and $y_{m}=-t$ results in an element in $\mathbb{Z}[x_{1},\dots,x_{n-1},y_{1},\dots,y_{m-1}]$, then we call $f(X,Y)$ a supersymmetric function. The elementary supersymmetric function and the complete supersymmetric function indexed by $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ are:
(2.6) $E_{\lambda}(X;Y)=\prod_{i=1}^{n}E_{\lambda_{i}}(X;Y),$
@@
$E_{r}(X;Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}e_{j}(X)h_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0,\end{cases}$
@@
(2.7) $H_{\lambda}(X;Y)=\prod_{i=1}^{n}H_{\lambda_{i}}(X;Y),$
@@
$H_{r}(X;Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}h_{j}(X)e_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0.\end{cases}$
It is convenient to define $E_{r}(X;Y)$ and $H_{r}(X;Y)$ to be zero for $r<0$.
@@
  • the entries increase weakly across rows and down columns,
  • the unprimed entries strictly increase down columns,
  • the primed entries strictly increase across rows.
@@
$\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{n_{i}(T)}{\prod_{j=1}^{m}y_{j}^{n_{j^{\prime}}(T)}},$
where $n_{\alpha}(T)$, for $\alpha$ being any of the entries given in 2.1, is the number of occurrences of $\alpha$ in $T$. Denote the set of supertableaux of skew shape $\lambda$ with entries given in 2.1 by $\operatorname{SSYT}_{n/m}(\lambda)$. For an integer partition $\lambda$, the hook Schur (or supersymmetric Schur) polynomial, denoted $\operatorname{hs}_{\lambda}(X;Y)$ is given by
(2.8) $\operatorname{hs}_{\lambda/\mu}(X;Y)\coloneqq\sum_{T\,\in\,\operatorname{SSYT}_{n/m}(\lambda)}\operatorname{wt}(T).$
@@
$\operatorname{hs}_{(2,1)}(x_{1},x_{2};y_{1})=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}y_{1}+2x_{1}x_{2}y_{1}+x_{2}^{2}y_{1}+x_{1}y_{1}^{2}+x_{2}y_{1}^{2}.$
The Jacobi–Trudi formula expresses $\operatorname{hs}_{\lambda}(X;Y)$ in terms of the complete supersymmetric functions [PT92].
(2.9) $\operatorname{hs}_{\lambda}(X;Y)=\det\left(H_{\lambda_{i}-i+j}(X;Y)\right)_{1\leq i,j\leq n}.$
Note that the hook Schur polynomial $\operatorname{hs}_{\lambda}(X;Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. Now we consider orthosymplectic Schur functions after defining symplectic Schur functions.
@@
A symplectic tableau $T$ of shape $\lambda$ is a filling of the shape with entries from $1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}$ satisfying the following conditions:
  • the entries increase weakly across rows and strictly down columns,
  • the entries in the $i^{\text{th}}$ row must be greater than or equal to $i$.
@@
$\operatorname{wt}(T)={\prod_{i=1}^{n}x_{i}^{n_{i}(T)-n_{\overline{i}}(T)},}$
where $n_{i}(T)$ and $n_{\overline{i}}(T)$ are the number of occurrences of $i$ and $\overline{i}$ in $T$ respectively. Denote the set of symplectic tableaux of shape $\lambda$ filled with entries given in 2.3 by $\operatorname{SP}_{n}(\lambda)$.
@@
The following is a symplectic tableau $T$ of shape $\lambda=(3,2,2)$ with $\operatorname{wt}(T)={x_{2}^{-1}}$.
@@
$\mathrm{sp}_{\lambda}(X)\coloneqq\sum_{T\in\operatorname{SP}_{n}(\lambda)}\operatorname{wt}(T),$
where the sum runs over the set of tableaux in $\operatorname{SP}_{n}(\lambda)$. The symplectic Schur function can be expressed using the Weyl character formula in the following way. Let $\lambda$ be a partition of length at most $n$. Then
(2.10) $\mathrm{sp}_{\lambda}(X)=\frac{\displaystyle\det\Bigr{(}x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}\Bigr{)}_{1\leq i,j\leq n}}{\displaystyle\det\Bigr{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}_{1\leq i,j\leq n}},$
@@
(2.11) $\det\Bigr{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}_{1\leq i,j\leq n}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\,\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
@@
An orthosymplectic tableau $T$ of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from $B$ satisfying the following conditions:
  • the portion $S$ of $T$ that contains entries only from $B_{0}$ is symplectic,
@@
$\operatorname{wt}(T)={\prod_{i=1}^{n}x_{i}^{n_{i}(T)-n_{\overline{i}}(T)}\prod_{j=1}^{m}y_{j}^{n_{i^{\prime}}(T)}},$
where $n_{\alpha}$ is the number of occurrences of $\alpha$, for $\alpha$ being any of the entries in $B$. Denote the set of orthosymplectic tableaux of shape $\lambda$ filled with entries in $B$ by $\operatorname{OSP}_{n/m}({\lambda})$.
@@
The following is an orthosymplectic tableau $T$ of shape $\lambda=(3,3,2,1)$ and $\operatorname{wt}(T)=x_{1}^{-1}x_{4}y_{1}y_{2}y_{3}^{3}$.
@@
$\operatorname{spo}_{\lambda}(X;Y)\coloneqq\sum_{T\,\in\,\operatorname{OSP}_{n/m}({\lambda})}\operatorname{wt}(T).$
Note that $\operatorname{spo}_{\lambda}(X;\emptyset)=\mathrm{sp}_{\lambda}(X)$. Moreover,
$\operatorname{spo}_{\lambda}(X;Y)=\sum_{\mu}\mathrm{sp}_{\mu}(X)\mathrm{s}_{\lambda^{\prime}/\mu^{\prime}}(Y).$
@@
Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then we have the following orthosymplectic tableaux:
@@
(2.13) $\operatorname{spo}_{\lambda}(X;Y)={\det\left(\begin{array}[]{c|c}\left(J_{\lambda_{i}-i+1}\right)_{1\leq i\leq n}&(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2})_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)},$
where $J_{r}=\sum_{l=0}^{r}h_{l}(X,\overline{X})e_{r-l}(Y)$. Here $h_{l}(X,\overline{X})$ is the Laurent polynomial obtained by substituting $x_{n+i}=\bar{x}_{i},1\leq i\leq n$ in $h_{l}({x_{1},x_{2},\dots,x_{2n}})$.
Note that the orthosymplectic Schur polynomial $\operatorname{spo}_{\lambda}(X;Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. The following remark gives an expression for $\operatorname{spo}_{\lambda}(X;y)$ if $\lambda_{n+1}\leq 1$.
@@
For $m=1$, suppose $\lambda$ is a partition such that $\ell(\lambda)=\ell$ and $\lambda_{n+1}\leq 1$. If $\ell\leq n$, then $\operatorname{spo}_{\lambda}(X;y)$ is given by 1.4. If $\ell>n$, then $\lambda_{n+1}=\cdots=\lambda_{\ell}=1$. So, by 2.5, $T$ contains $1^{\prime}$ in all the boxes of $\lambda$ from the $(n+1)^{\text{th}}$ row to $\ell^{\text{th}}$ row, for all $T\in\operatorname{OSP}_{n/1}(\lambda)$. Therefore, using (2.10) and (2.11) we have
$\operatorname{spo}_{\lambda}(X;y)=\frac{y^{\ell-n}}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})\right),$
@@
Let $\lambda$ be a partition of length at most $n$. An $(n-1,1)$-symplectic tableau of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from
@@
  • the entries increase weakly across rows and strictly down columns,
  • the entries of the $i^{\text{th}}$ row are greater than or equal to $i$.
We denote by ${\operatorname{SP}^{(n-1,1)}(\lambda)}$ the set of $(n-1,1)$-symplectic tableaux of shape $\lambda$. Given an $(n-1,1)$-symplectic tableau $T$, we define
@@
where $\alpha_{i}(T)$, $\alpha_{\overline{i}}(T)$ for $i\in[n-1]$, and $\alpha_{n}(T)$ are the number of occurences of $i$, $\overline{i}$ and $n$ respectively.